OpenStudy (anonymous):

if angle AOC = 85 degrees, angle BOC = 2x+10, and angle AOB =4x-15, find the degree measures of angle BOC and angle AOB.

OpenStudy (anonymous):

@stoopkid

OpenStudy (anonymous):

Lol ok: So you have one big angle AOB, and this is split into two angles by a line/point C, so you have two angles joined together: AOC and BOC, and the sum of these angles will be your total angle AOB. AOC + BOC = AOB. 85 + 2x + 10 = 4x - 15. Using algebra: 110 = 2x. x = 55. Substituting this value back into our equations: BOC = 2*55 + 10 = 110+10 = 120 AOB = 4*55 -15 = 220 - 15 = 205. Let's make sure this makes sense: AOC + BOC = 85 + 120 = 205 = AOB. So it works out!

jhonyy9 (jhonyy9):



OpenStudy (anonymous):

so BOC is 55

jhonyy9 (jhonyy9):

hope this will help you understanding easy

OpenStudy (anonymous):

if BOC is 55 what is AOB?

OpenStudy (anonymous):

BOC is not 55. x is 55. BOC is 2x + 10 = 120. :)

OpenStudy (anonymous):

i guess you need to take AOB+BOC=AOC if AOC is considered to be the largest angle

OpenStudy (anonymous):

in this fashion we get the angles as 45 and 40 there can be multiple solutions unless the figure is specified

OpenStudy (anonymous):

Yes, my answer is only valid if the largest angle is AOB and C is somewhere in between the two.

OpenStudy (anonymous):

a. BOC=30 and AOB=55 b. BOC=40 and AOB=45 c. BOC=45 and AOB=40 d. BOC=55 and AOB=30 @stoopkid

OpenStudy (anonymous):

Ok, giligails solution must be the correct one

OpenStudy (anonymous):

the 2 angles must either add up to 85 or (360-85)

OpenStudy (anonymous):

what seance mean?